DESIGN AND PERFORMANCE OF A DIRECT DIGITAL CHIRPED-PULSE FOURIER TRANSFORM MI-CROWAVE (CP-FTMW) SPECTROMETER OPERATING IN THE 2 - 8 GHZ FREQUENCY RANGE

\underline{STEVEN T. SHIPMAN}, LEONARDO ALVAREZ-VALTIERRA, JUSTIN L. NEILL, and BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904; ALBERTO LESSARI, Universidad de Valladolid, Departmento Química Física y Química Inorgánica, Facultad de Ciencias, Prado de la Magdalena, s/n, 47005 Valladolid, Spain; Z. KISIEL, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warszawa, Poland.

We have constructed a chirped-pulse Fourier transform microwave (CP-FTMW) spectrometer operating in the 2 - 8 GHz frequency range. The chirped excitation pulse executes a linear frequency sweep covering 2 - 8 GHz in pulses with duration of 9 μs or less. This chirped pulse is created using the 8-bit D/A converter with 20 Gs/s sample rate in a high-speed arbitrary waveform generator (AWG). The output of the AWG is sent to a high-power pulsed microwave amplifier. We have used both a 300W traveling wave tube amplifier and a 4W solid state amplifier in the spectrometer. The amplified chirped pulse is broadcast onto the molecular beam sample using a WRD250 double-ridge standard gain horn with 15 dBi gain. A second WRD250 gain horn is used to collect the broadband rotational free induction decay. The molecular emission is amplified by a broadband, low-noise amplifier and converted back to a digital signal using the 8-bit A/D converter with 20 Gs/s sample rate in a digital oscilloscope. The frequency domain spectrum is obtained through fast Fourier transform of the time-domain averaged FID with 40 μs duration. The performance of the spectrometer is demonstrated by measurment of the pure rotational spectrum of iodobenzene, its ^{13}C isotopomers, and a weakly bound complex with the neon carrier gas in the molecular beam expansion. A frequency accuracy of 4 kHz for the spectrometer is demonstrated by comparison to previous measurements of the iodobenzene spectrum using several spectrometers.^{a} Accurate transition intensities are obtained for both number density (^{13}C isotopomer signals) and transition moment (nuclear quadrupole hyperfine patterns). The open interaction region between the horn antennas makes it easy to incorporate a “Stark cage” for dipole moment determination.^{b} Performance of the broadband dipole moment determination of iodobenzene is tested by comparison to previous work. The implementation of broadband microave-microwave double-resonance spectroscopy is also presented for the iodobenzene spectrum.

aDorosh, O. et al., J. Mol. Spec. 246 (2007), 228-232.

bEmilsson, T. et al., J. Chem. Phys. 112 (2000), 1287-1294.